OpenStudy (anonymous):

f g(x) = 3x + 5, what is g–1(x)?

OpenStudy (carniel):

3x+5= -1 SO that means x=6

OpenStudy (carniel):

If im correct it should :)

OpenStudy (anonymous):

g^-1 if any one finds that confusing

OpenStudy (anonymous):

if? \[g^{-1}(x)=\frac{x-5}{3}\] original function says "multiply by 3, add 5" inverse says "subtract 5, divide by 3"

OpenStudy (anonymous):

like if you were solving \[3x+5=17\] you would subtract 5 and get \[3x=12\] then divide by 3 and get \[x=4\] that is what the inverse function does, take you back to where you started from

OpenStudy (anonymous):

ok , well thats what i got but the way ur explaining it confuses me

OpenStudy (anonymous):

g(x)let=y y=3x+5 x=3y-5 3y=x-5 3y/3=x/3-5/3 y=x/3-5/3 g(x)=x/3-5/3